### v4

# SL(2N,C) Yang-Mills Theories –

Abstract
We develop a four-dimensional gauge-gravity unification based on the $SL(2N,C)$ gauge theory taken in a universal Yang–Mills type setting. The accompanying tetrads are promoted to dynamical fields whose contracted invertibility condition is interpreted as a nonlinear sigma-model type length constraint. This triggers tetrad condensation and spontaneously breaks $SL(2N,C)\rightarrow SL(2,C)\times SU(N)$, lifting all noncompact directions. A special ghost-free curvature-squared Lagrangian provides a consistent quadratic sector, while an Einstein–Cartan linear curvature term is induced radiatively from fermion loops. Below the breaking scale, only a neutral tetrad associated with graviton and $SU(N)$ vector fields remain massless, whereas axial-vector and tensor fields of the entire gauge multiplet acquire heavy masses. The matter sector clearly points to a deeper elementarity of $SL(2N,C)$ spinors, which can be identified with preon constituents whose bound states form the observed quarks and leptons. Anomaly matching between preons and composites singles out $N=8$. The chain $SL(16,C)\rightarrow SL(2,C)\times SU(8)$ then naturally yields three composite quark–lepton families, while filtering out extraneous heavy states.
1 Introduction
It has long been recognized [1, 2, 3] that gravity can be cast as a gauge theory in close analogy with the gauge description of electromagnetic, weak, and strong interactions. In this language the spin connection associated with local Lorentz transformations plays the role of a gauge field, much as the Yang–Mills bosons do for internal symmetry groups. Once this parallel is taken seriously, it is natural to ask whether both gravitational and internal gauge symmetries can be embedded into a single non-compact group so that all interactions are described within a common gauge-theoretic framework that has been intensively discussed in various forms [4, 5, 6, 7, 8, 9, 10]. However, despite an attractive algebraic structure, these schemes face well-known obstacles which we consider below in the unified $SL(2N,C)$ symmetry case.
For the gravitational sector of such unification the relevant local symmetry is $SL(2,C)$, the double cover of the proper orthochronous Lorentz group, which furnishes the appropriate spinorial representation for fermions in four-dimensional spacetime. If one assumes that the fundamental spinor fields are also responsible for organizing the internal quantum numbers, it is then tempting to extend $SL(2,C)$ to a larger group $SL(2N,C)$ in which $SU(N)$ appears as an internal symmetry subgroup. In the following we refer to such $SL(2N,C)$ based constructions as hyperunified theories (HUTs), since they collect together $SL(2,C)$ gauge gravity and $SU(N)$ grand unification within a single local group. The internal $SU(N)$ piece will be called hyperflavor, in order to distinguish it from the Lorentzian $SL(2,C)$ sector which remains neutral under internal charges. The full $SL(2N,C)$ gauge multiplet includes spin-$1$ (vector and axial-vector) and spin-$2$ (bi-vector or tensor) fields, which in principle may unify Standard Model interactions and gravity—pending resolution of known challenges.
First, the unavoidable axial-vector and tensor partners of the usual vector fields have no experimental support and therefore must either decouple or lie at very high energies. Moreover, while internal forces are naturally described by Yang–Mills kinetic terms quadratic in the field strength, gravity in its Einstein or Einstein–Cartan form is linear in curvature, so a naive unified Lagrangian typically contains both linear and quadratic terms with different couplings. This tends to spoil the idea of a fully unified gauge sector and generically yields multiple tensor “gravitons” carrying $SU(N)$ quantum numbers, together with ghost or tachyon modes in standard $R+R^{2}$ constructions.
A further difficulty is that the matter sector of a hyperunified $SL(2N,C)$ theory is strongly constrained. Any non-fundamental representation of $SL(2N,C)$ unavoidably mixes fields of different spins inside a single multiplet and therefore cannot describe purely fermionic states such as quarks and leptons. Conversely, the fundamental $2N$-dimensional representation contains only spin-$\tfrac{1}{2}$ fields, but its internal $SU(N)$ quantum numbers do not match the observed pattern of Standard-Model charges. Thus, the $SL(2N,C)$ cannot act directly on quarks and leptons without introducing a huge number of unobserved higher-spin or exotic-charged states. This strongly suggests that quarks and leptons might not be truly elementary in the $SL(2N,C)$ sense, but instead emerge as composites of more fundamental preons. If these preons are chiral spin-$1/2$ fermions transforming in the fundamental representation of $SL(2N,C)$, the group becomes a natural candidate for organizing both gravitational and internal dynamics at very short distances.
Finally, a conceptual issue concerns the Coleman–Mandula theorem [11], which severely restricts nontrivial unifications of spacetime and internal symmetries within a simple Lie algebra at the S-matrix level. As we shall see, the mechanism we propose circumvents the obvious conflict by effectively lifting the non-compact directions through tetrad dynamics, so that the residual low-energy symmetry is of standard product form.
The central idea in this work is that the tetrad, when treated as a genuine dynamical field, can play three simultaneous roles. First, its invertibility (orthonormality) condition can be interpreted as a nonlinear $\sigma$-model type length constraint which forces a nonzero vacuum expectation value. Second, this VEV spontaneously breaks $SL(2N,C)$ down to $SL(2,C)\times SU(N)$, such that only the $SU(N)$ vector fields remain light while the axial and tensor components become heavy. Third, in the presence of suitable fermion multiplets the same tetrad and tensor fields generate an Einstein–Cartan linear curvature term radiatively, even if such a term is absent at tree level.
To make the quadratic-curvature sector consistent we adopt a particular ghost-free combination of curvature-squared terms originally proposed by Neville [12, 13]. We show that this structure arises naturally from an $SL(2N,C)$ invariant quadratic Lagrangian once the tetrad is restricted to be neutral in hyperflavor space. With this choice the strength-squared Lagrangian becomes fully unified: a single gauge coupling governs the dynamics of the vector, axial, and tensor submultiplets.
Matter fields in $SL(2N,C)$ are then reconsidered from a composite perspective. We argue that a preon model with chiral fermions in fundamental $SL(2N,C)$ representations, together with metacolor confinement and anomaly matching, picks out $SL(16,C)$ with an $SU(8)$ hyperflavor subgroup as a particularly compelling case. In this scenario three chiral quark-lepton families arise as composites of preons and fit naturally within a single $SU(8)$ based structure.
The paper is organized as follows. Section 2 reviews the construction of $SL(2N,C)$ gauge theories, beginning with neutral spinors and $SL(2,C)$ symmetry, and then extending to charged spinors and full $SL(2N,C)$ hyperunification. Section 3 develops the quadratic strength-tensor Lagrangians, identifying a ghost-free choice that treats all gauge submultiplets on equal footing. In Section 4 we introduce dynamical tetrads, implement the nonlinear length constraint, and analyze the spontaneous breaking $SL(2N,C)\rightarrow SL(2,C)\times SU(N)$, including the Coleman–Weinberg selection of a hyperflavor-blind vacuum and the decoupling of hyperflavored tetrad modes. Section 5 explains how an Einstein–Cartan gravity term is induced at one loop by fermion bubbles involving tetrads and tensor fields. Section 6 turns to the matter sector and shows how preon-based $SL(16,C)$ hyperunification can accommodate three composite quark-lepton families. Section 7 summarizes the main results and outlines directions for further work.
2 SL(2N,C)\ SL(2N,C) gauge theories
Assuming that spinor fields provide the fundamental building blocks of both spacetime and internal structure leads naturally from the $SL(2,C)$ gauge group of gravity [3] to an enlarged local symmetry $SL(2N,C)$, where $N$ measures the size of an internal $SU(N)$ subgroup. This $SU(N)$ is interpreted as a hyperflavor symmetry unifying all known gauge quantum numbers of quarks and leptons—color, weak isospin, and family indices included. In what follows, we outline how such a hyperflavor symmetry arises and how it feeds into hyperunification.
2.1 Neutral spinors – $SL(2,C)$ symmetry
We start with the $SL(2,C)$ gauge symmetry case for a neutral spinor. Consider a Dirac field $\Psi$ which transforms in the fundamental spinor representation of $SL(2,C)$, acting in the tangent (spinor) bundle at each spacetime point:
$\Psi\rightarrow\Omega\Psi\text{, \ \ }\Omega=\exp\left\{\frac{i}{4}\theta_{ab}\gamma^{ab}\right\}\text{ }$ (1)
where the matrix $\Omega$ satisfies the pseudounitarity condition
$\Omega^{-1}=\gamma_{0}\Omega^{+}\gamma_{0}\,.$ (2)
To preserve the invariance of the free kinetic term $i\overline{\Psi}\gamma^{\mu}\partial_{\mu}\Psi$ under local $SL(2,C)$ transformations, the constant matrices $\gamma^{\mu}$ must be replaced by tetrad-valued matrices $e^{\mu}$ that transform as
$e^{\mu}\rightarrow\Omega e^{\mu}\Omega^{-1}\text{ }$ (3)
and are related to the usual tetrad fields $e^{a}{}_{\mu}$ by
[…]
Infinitesimally, this implies for the tetrad components
$\delta e^{\mu c}=\frac{1}{2}\theta_{ab}(e^{\mu a}\eta^{bc}-e^{\mu b}\eta^{ac})$ (5)
while the usual invertibility conditions
$e_{\mu}^{a}e_{a}^{\nu}=\delta_{\mu}^{\nu}\text{, \ }e_{\mu}^{a}e_{b}^{\mu}=\delta_{b}^{a}$ (6)
ensure that these tetrads define a nondegenerate metric,
[…]
Promoting $SL(2,C)$ to a local symmetry, $\theta_{ab}\rightarrow\theta_{ab}(x)$, requires introducing a spin-connection one-form $I_{\mu}$ transforming as
$I_{\mu}\rightarrow\Omega I_{\mu}\Omega^{-1}-\frac{1}{ig}(\partial_{\mu}\Omega)\Omega^{-1}$ (8)
and defining the covariant derivative
[…]
with gauge coupling $g$. The spin connection can be expanded in terms of Dirac generators as
$I_{\mu}=\frac{1}{4}T_{\mu[ab]}\gamma^{ab}$ (10)
where $T_{\mu}{}^{[ab]}$ transforms as
$\delta T_{\mu}^{[ab]}=\frac{1}{2}\theta_{[cd]}[(T_{\mu}^{[ac]}\eta^{bd}-T_{\mu}^{[ad]}\eta^{bc})-(T_{\mu}^{[bc]}\eta^{ad}-T_{\mu}^{[bd]}\eta^{ac})]-\frac{1}{g}\partial_{\mu}\theta^{[ab]}$ (11)
i.e. as a Lorentz-antisymmetric tensor in the local frame and a spacetime one-form.
Its field strength
$T_{\mu\nu}^{[ab]}=\partial_{[\nu}T_{\mu]}^{[ab]}+g\eta_{cd}T_{[\mu}^{[ac]}T_{\nu]}^{[bd]}\,$, (12)
when appropriately squared (being contracted with tetrad as well), leads to an $SL(2,C)$ invariant Yang-Mills type Lagrangian. However, in the presence of a tetrad, an additional invariant term linear in $T_{\mu\nu}^{[ab]}$, can also be written as
$e\mathcal{L}_{G}^{(1)}=\frac{1}{2\kappa}e_{[a}^{\mu}e_{b]}^{\nu}T_{\mu\nu}^{[ab]}\text{ , \ \ }e\equiv[-\det\mathrm{Tr}(e^{\mu}e^{\nu})/4]^{-1/2}$ (13)
where $\kappa$ is a dimensionful coefficient. Identifying $\kappa$ with $8\pi/M_{Pl}^{2}$ yields the Einstein–Cartan action in Palatini formulation [3, 14]. Varying with respect to $T_{\mu}^{ab}$ expresses it algebraically in terms of the tetrad and its derivatives, and one recovers the usual Einstein action plus torsion-induced four-fermion terms. The factor $e$ ensures invariance under general four-dimensional coordinate transformations, $GL(4,R)$ as well [3]. The underlying spacetime is still globally Minkowskian; the $SL(2,C)$ gauge group acts on the tangent (spinor) space rather than on coordinates.
The gauge coupling the fermion to $I_{\mu}$ via (9, 10) produces an additional interaction term
$e\mathcal{L}_{M}=-\frac{1}{2}g\epsilon^{abcd}T_{\mu[ab]}e_{c}^{\mu}\overline{\Psi}\gamma_{d}\gamma_{5}\Psi$ (14)
which, after eliminating torsion, leads to an effective four-fermion contact term
$\kappa\left(\overline{\Psi}\gamma_{d}\gamma^{5}\Psi\right)(\overline{\Psi}\gamma^{d}\gamma^{5}\Psi)$ (15)
characteristic of Einstein–Cartan gravity [2].
We show below, in Sec. 5, that in the $SL(2N,C)$ hyperunified theory, whose fundamental bosonic sector is governed by a universal quadratic-strength Lagrangian with a single gauge coupling, the Einstein–Cartan term need not be introduced as an independent tree-level input: it can instead be generated radiatively by the corresponding fermion-loop diagrams.
2.2 Charged spinors – $SL(2N,C)$ hyperunification
We now pass to the full $SL(2N,C)$ group, which contains the spacetime symmetry $SL(2,C)$ and an internal hyperflavor $SU(N)$ among its primary subgroups. The $(8N^{2}-2)$ generators of $SL(2N,C)$ can be written as tensor products of Dirac matrices with $SU(N)$ generators. A convenient parametrization [4] of a general transformation acting on fermions is
$\Omega=\exp\left\{\frac{i}{2}\left[\left(\theta^{k}+\theta_{5}^{k}\gamma_{5}\right)\lambda^{k}+\frac{1}{2}\theta_{ab}^{K}\gamma^{ab}\lambda^{K}\right]\right\}\text{ \ \ }(K=0,k)$ (16)
where $\lambda^{k}$ ($k=1,\dots,N^{2}-1$) are the standard $SU(N)$ Gell-Mann matrices and $\lambda^{0}=\sqrt{2/N}\,\mathbf{1}_{N}$ denotes the $U(1)$ generator. The parameters $\theta^{k}$, $\theta_{5}^{k}$ and $\theta_{ab}^{K}$ may be global or local [^1] The $\gamma$ and $\lambda$ matrices obey $\gamma^{ab}$ $=$ $i[\gamma^{a},\gamma^{b}]/2,\text{ }\{\gamma^{a},\gamma^{b}\}=2\eta^{ab}\mathbf{1},\text{ }$ $[\lambda^{k},\lambda^{l}]$ $=$ $2if^{klm}\lambda^{m},\text{ }\{\lambda^{k},\lambda^{l}\}=\frac{4}{N}\delta^{kl}\widehat{1}+2d^{klm}\lambda^{m}\,.$ . Uppercase indices $I,J,K$ will refer to $U(1)\times SU(N)$ components, while lowercase $i,j,k$ will be used for the purely $SU(N)$ part. The equation (16) is written in a Dirac-matrix form because we are describing the action of the $SL(2N,C)$ symmetry on the fermion multiplet $\Psi$, i.e. on Dirac spinors carrying an additional hyperflavor $SU(N)$ index (so $\Psi$ lives in the space $\mathbb{C}^{4}\times$ $\mathbb{C}^{N}$). In this Dirac $\times$ hyperflavor realization the transformation is naturally expressed using $\gamma$ matrices; in particular, in a chiral basis it is reducible and acts separately on left- and right-handed components. This is completely analogous to the pure $SL(2,C)$ case discussed above in (1) for a neutral Dirac fermion living in the space $\mathbb{C}^{4}$.
The tetrad multiplet in the full $SL(2N,C)$ theory generically has the form
$e_{\mu}=(e_{\mu}^{aK}\gamma_{a}+e_{\mu 5}^{aK}\gamma_{a}\gamma_{5})\lambda^{K}$ (17)
and transforms as in (3), now with $\Omega$ given by (16). One can simplify this structure by eliminating the axial component through the introduction of an auxiliary scalar multiplet
$S=\exp\{i[(s^{k}+p^{k}\gamma_{5})\lambda^{k}+t_{ab}^{K}\gamma^{ab}\lambda^{K}/2]\}$ (18)
transforming as $S\rightarrow\Omega S$. The combination $S^{-1}eS$ can be arranged, by a suitable choice of background values for $S$, to remove the axial-vector part of the tetrad [4]. After this reduction, we adopt the purely vectorial tetrad form
$e_{\mu}=e_{\mu}^{aK}\gamma_{a}$ (19)
for the rest of the analysis.
2.3 Gauging $SL(2N,C)$
Localizing the transformation (16) introduces a gauge field $I_{\mu}$ transforming as
$I_{\mu}\rightarrow\Omega I_{\mu}\Omega^{-1}-\frac{1}{ig}(\partial_{\mu}\Omega)\Omega^{-1}$ (20)
with field strength
$I_{\mu\nu}=\partial_{[\mu}I_{\nu]}+ig[I_{\mu},I_{\nu}]$ (21)
and fermionic covariant derivative
$\partial_{\mu}\Psi\rightarrow D_{\mu}\Psi=\partial_{\mu}\Psi+igI_{\mu}\Psi\,.$ (22)
The gauge connection $I_{\mu}$ can be decomposed into vector, axial, and tensor parts:
$I_{\mu}=V_{\mu}+A_{\mu}+T_{\mu}=\frac{1}{2}\left(V_{\mu}^{k}+A_{\mu}^{k}\gamma_{5}\right)\lambda^{k}+\frac{1}{4}T_{\mu[ab]}^{K}\gamma^{ab}\lambda^{K}\text{ \ \ }(K=0,k)$ (23)
where each submultiplet carries a spacetime index $\mu$ and appropriate internal indices. [^2] All three submultiplets transform as four-vectors in spacetime. In the local Lorentz frame $V$ and $A$ behave as scalar and pseudoscalar fields, respectively, while $T$ transforms as an antisymmetric rank-2 tensor. The corresponding contribution to the field strength reads
$I_{\mu\nu}$ $=$ $\frac{1}{2}\partial_{[\mu}\left(V^{k}+A^{k}\gamma_{5}\right)_{\nu]}\lambda^{k}-\frac{1}{2}f^{ijk}g\left(V^{i}+A^{i}\gamma_{5}\right)_{\mu}(V^{j}+A^{j}\gamma_{5})_{\nu}\lambda^{k}$ (24)
$+\frac{1}{4}\left(\partial_{[\mu}T_{\nu]}^{[ab]K}\gamma_{ab}\lambda^{K}+i\frac{g}{4}T_{\mu}^{[ab]K}T_{\nu}^{[a^{\prime}b^{\prime}]K^{\prime}}[\lambda^{K}\gamma_{ab},\lambda^{K^{\prime}}\gamma_{a^{\prime}b^{\prime}}]\right)$
and the generic gauge-invariant matter coupling can be written in terms of $e^{\mu}$ and the submultiplets as
$e\mathcal{L}_{M}=-\frac{g}{2}\overline{\Psi}\left\{e^{\mu},\left[\frac{1}{2}\left(V_{\mu}^{k}+A_{\mu}^{k}\gamma_{5}\right)\lambda^{k}+\frac{1}{4}T_{\mu[ab]}^{K}\gamma^{ab}\lambda^{K}\right]\right\}\Psi\,.$ (25)
Thus all three types of gauge fields couple with the same universal coupling $g$ in the underlying $SL(2N,C)$ theory, which typically includes a standard quadratic field-strength term for each submultiplet, while a possible linear curvature term, analogous to (13), may induce radiatively (Sec. 5). The key question is how the observed low-energy spectrum can retain only the $SU(N)$ vector fields and a single graviton, while the axial-vector and tensor states are decoupled.
As will be seen below, a crucial simplification occurs once tetrads are constrained to be neutral in hyperflavor space. In that case only the $V_{\mu}^{k}$ and the singlet tetrad $e_{\mu}^{a0}$ remain relevant for low-energy physics: the former mediate internal forces, while the latter underlies the induced Einstein–Cartan gravity.
2.4 Tetrads linked to gravity
The general tetrad ansatz (19) includes components carrying hyperflavor indices. However, the invertibility conditions suitable for coupling to a four-dimensional metric severely restrict this structure. Let us assume
$e_{\mu}=e_{\mu}^{aK}\gamma_{a}\lambda^{K},\text{ }e_{\mu}^{aK}e_{b}^{\mu K^{\prime}}=\Delta_{b}^{aKK^{\prime}},\text{ }e_{\mu}^{aK}e_{a}^{\nu K^{\prime}}=\Delta_{\mu}^{\nu KK^{\prime}}$ (26)
with some matrices $\Delta$ whose form is to be determined. In the pure gravity limit, we expect
$\Delta_{b}^{a00}=\delta_{b}^{a}\text{ , \ }\Delta_{\mu}^{\nu 00}=\delta_{\mu}^{\nu}\,.$ (27)
Multiplying (26) by $e_{\rho}^{bK^{\prime\prime}}$ and $e_{a}^{\rho K^{\prime\prime}}$ and using basic algebraic identities one finds
$e_{\mu}^{aK}=\Delta_{b}^{aK0}e_{\mu}^{b0}\text{, \ }e_{a}^{\mu K^{\prime}}=\Delta_{a}^{bK^{\prime}0}e_{b}^{\mu 0}$ (28)
which leads to
$\Delta_{b}^{aKK^{\prime}}=\Delta_{c}^{aK0}\Delta_{b}^{cK^{\prime}0}$ (29)
and similarly
$\Delta_{\mu}^{\nu KK^{\prime}}=\Delta_{\rho}^{\nu K0}\Delta_{\mu}^{\rho K^{\prime}0}\,.$ (30)
Assuming constant multiplicative factors for $\Delta$, one finds that a consistent solution compatible with (27) is
$e_{\mu}^{aK}e_{b}^{\mu K^{\prime}}=\Delta_{b}^{aKK^{\prime}}=\delta_{b}^{a}\delta^{K0}\delta^{K^{\prime}0}\text{ , \ }e_{\mu}^{aK}e_{a}^{\nu K^{\prime}}=\Delta_{\mu}^{\nu KK^{\prime}}=\delta_{\mu}^{\nu}\delta^{K0}\delta^{K^{\prime}0}$ (31)
i.e. the standard invertibility relations hold only for the $K=0$ component. This implies that the tetrad must effectively lie in the $SL(2,C)$ subgroup:
$e_{\mu}^{aK}=e_{\mu}^{a}\delta^{K0}$ (32)
if it is to define a conventional spacetime metric. In other words, only a single $SU(N)$-singlet tetrad is compatible with the usual tangent bundle structure.
One can enforce this by imposing (32) as a background constraint. Alternatively, when tetrads are treated as dynamical fields, neutral tetrads may emerge via the spontaneous breaking of $SL(2N,C)$. This way, inspired by the nonlinear $\sigma$-model type length constraint
$\frac{1}{4}e_{\mu}^{aK}e_{a}^{\mu K}=1$ (33)
which, when applied to a general tetrad, drives spontaneous symmetry breaking down to $SL(2,C)\times SU(N)$. For such a vacuum configuration, tetrads written as an appropriate sum of symmetric and antisymmetric parts,
$e_{\mu}^{aK}\,=\eta_{\mu b}\left[\,\eta^{ab}\delta^{K0}+\widetilde{e}^{\{ab\}K}+\widehat{e}^{[ab]K}\right]$ (34)
should provide an extremum of the action with $\delta_{\mu}^{a}\delta^{K0}$ acting as the effective Higgs mode, while $\eta_{\mu b}\widehat{e}^{[ab]K}$ and $\eta_{\mu b}\widetilde{e}^{\{ab\}k}$ are zero modes. Absorption of these modes induces, as we see later in Sec. 4, masses for the submultiplets of tensor and axial-vector fields, $T_{\mu}^{[ab]K}$ and $A_{\mu}^{k}$, respectively, which are associated with the broken coset generators beyond the residual $SL(2,C)\times SU(N)$ symmetry.
3 Lagrangians
An $SL(2N,C)$-invariant action may contain both linear and quadratic terms in the field strengths, each with its own coupling. For our purposes we focus on the quadratic (Yang–Mills type) sector and require it to be governed by a single universal gauge coupling, common to vector, axial, and tensor parts of the connection (23). As will be seen in the next section, the Einstein–Cartan linear curvature term, while not excluded at tree level by symmetry, may be viewed as an emergent contribution generated primarily by fermion loops. In this way one can maintain hyperunification in the quadratic sector while letting gravity arise from quantum effects.
3.1 Vector and axial-vector fields
We first examine the part of the Lagrangian associated with the $SU(N)$ vector and axial-vector fields, which are the basic spin-1 carriers of hyperflavor. From the quadratic strength term built out of (24) one finds
$e\mathcal{L}_{H}^{VA}$ $=$ $-\frac{1}{2}\mathrm{Tr}[(V_{\mu\nu}^{k}\lambda^{k}+\gamma_{5}A_{\mu\nu}^{k}\lambda^{k})^{2}/4]=-\frac{1}{4}(V_{\mu\nu}^{k})^{2}-\frac{1}{4}(A_{\mu\nu}^{k})^{2}$
$=$ $-\frac{1}{4}[\partial_{[\mu}V_{\nu]}^{k}-gf^{ijk}(V_{\mu}^{i}V_{\nu}^{j}+A_{\mu}^{i}A_{\nu}^{j})]^{2}-\frac{1}{4}[\partial_{[\mu}A_{\nu]}^{k}-gf^{ijk}(V_{\mu}^{i}A_{\nu}^{j})]^{2}$
where $V_{\mu\nu}^{k}$ and $A_{\mu\nu}^{k}$ are the corresponding field strengths. The vector fields thus have the standard $SU(N)$ gauge structure, while the axial vectors behave as adjoint matter with self-interactions induced by the coupling to $V_{\mu}^{k}$. Some interactions involving $A_{\mu}^{k}$ and $T_{\mu}$ break the pure $SU(N)$ gauge invariance, but these terms will be irrelevant at low energies once tensor fields decouple.
In the matter sector, the vector fields couple to ordinary fermions through
$e\mathcal{L}_{HM}^{V}=-\frac{g}{2}V_{\mu}^{i}\overline{\Psi}e_{a}^{\mu}\gamma^{a}\lambda^{i}\Psi$ (36)
whereas the axial-vector couplings vanish for a neutral tetrad, so $A_{\mu}^{k}$ have no direct interactions with fermionic matter. They can still participate in loop processes through their self-interactions with vectors. Present experimental data [15] do not show direct evidence for axial gauge bosons, but do not completely forbid heavy states with such quantum numbers in the multi-TeV range. Alternatively, axial vectors might develop a condensate $\langle A_{\mu}^{k}\rangle=\mathfrak{n}_{\mu}^{k}\mathcal{M}$ at very high scales, where $\mathfrak{n}_{\mu}^{k}$ is a unit Lorentz vector, $\mathfrak{n}_{k}^{\mu}\mathfrak{n}_{\mu}^{k}=1$. In such a background the vector sector recovers gauge invariance, while Lorentz symmetry is spontaneously—and possibly only very weakly—broken [10].
3.2 Total gauge multiplet
As discussed earlier, a consistent hyperunified $SL(2N,C)$ theory should allow curvature-squared invariants for the tensor sector which are free of ghosts and tachyons, while remaining compatible with the quadratic terms for vectors and axial vectors. Among the candidate models, the ghost-free combination identified by Neville [12] (and further analyzed in [13]) stands out as particularly suitable. In the pure $SL(2,C)$ case, the corresponding Lagrangian reads
$e\mathcal{L}_{G}^{(2)}=\lambda T_{abcd}\left(T^{abcd}-4T^{acbd}+T^{cdab}\right)$ (37)
where $T^{abcd}=T_{\mu\nu}^{ab}e^{\mu c}e^{\nu d}$ and $T_{\mu\nu}{}^{ab}$ is the curvature of the spin connection. Writing everything in terms of the tensor fields and tetrads gives
$e\mathcal{L}_{G}^{(2)}=\lambda(T_{ab}^{\mu\nu}T_{\mu\nu}^{ab}-4T_{ab}^{\mu\nu}T_{\rho\nu}^{ac}e_{\mu c}e^{\rho b}+T_{ab}^{\mu\nu}T_{\rho\sigma}^{cd}e_{\mu c}e_{\nu d}e^{\rho a}e^{\sigma b})$ (38)
where $\lambda$ is a constant. In this model the spin connection becomes propagating, with a massive excitation in addition to the massless graviton. For generic parameters the extra mode has a mass $\sim M_{P}^{2}/\lambda$ and is expected to be phenomenologically irrelevant unless $\lambda$ is very large.
We now embed this structure into the hyperunified $SL(2N,C)$ theory by writing a general quadratic Lagrangian
$e\mathcal{L}_{H}^{(2)}=\lambda_{I}\mathrm{Tr}(aI^{\mu\nu}I_{\mu\nu}+bI^{\mu\nu}I_{\rho\nu}e_{\mu}e^{\rho}+cI^{\mu\nu}I_{\rho\sigma}e_{\mu}e_{\nu}e^{\rho}e^{\sigma})$ (39)
with constants $a,b,c$ to be fixed by the requirement that the $SL(2,C)$ limit of (39) reproduces (38). Using neutral tetrads satisfying (32) and the trace properties of Dirac and $SU(N)$ matrices, one finds for the vector and axial-vector contributions
$\mathrm{Tr}\left(I^{(W)\mu\nu}I_{\mu\nu}^{(W)}\right)$ $=$ $(W_{\mu\nu}^{k}W^{\mu\nu k})/2,\text{ }$
$\mathrm{Tr}\left(I^{(W)\mu\nu}I_{\rho\nu}^{(W)}e_{\mu}e^{\rho}\right)$ $=$ $2(W_{\mu\nu}^{k}W^{\mu\nu k}),$
$\mathrm{Tr}\left(I^{(W)\mu\nu}I_{\rho\sigma}^{(W)}e_{\mu}e_{\nu}e^{\rho}e^{\sigma}\right)$ $=$ $4(W_{\mu\nu}^{k}W^{\mu\nu k})\text{ \ }(W\equiv V,A)$ (40)
and for the tensor submultiplet
$\mathrm{Tr}\left(I^{(T)\mu\nu}I_{\mu\nu}^{(T)}\right)$ $=$ $T_{ab}^{\mu\nu K}T_{\mu\nu}^{abK},\text{ }$
$\mathrm{Tr}\left(I^{(T)\mu\nu}I_{\rho\nu}^{(T)}e_{\mu}e^{\rho}\right)$ $=$ $3T_{ab}^{\mu\nu K}T_{\mu\nu}^{abK}+4(T_{ab}^{\mu\nu K}T_{\rho\nu}^{acK}e_{\mu c}e^{\rho b}),$
[…]
Combining these pieces, the quadratic Lagrangian involves
$(a/2+3b+4c)W^{2}+(a+3b+2c)TT+4b\,TTee+8c\,TTeeee$ (42)
where $W^{2}$ denotes the vector/axial-vector term and the symbols $TT$ etc. represent the tensor structures in (38).
To reproduce the pure-gravity Neville combination (38) one must impose
$a+3b+2c=1,\text{ }b=-1,\text{ }c=1/8\rightarrow a=15/4$ (43)
which fixes $a,b,c$ and yields a coefficient $(3/8)$ for $W^{2}$ in (42). With these values the hyperunified quadratic Lagrangian becomes
$e\mathcal{L}_{H}^{(2)}$ $=$ $-\frac{1}{4}(V_{\mu\nu}^{k}V^{\mu\nu k}+A_{\mu\nu}^{k}A^{\mu\nu k})$
$-\frac{1}{4}\left(T_{ab}^{\mu\nu K}T_{\mu\nu}^{abK}-4T_{ab}^{\mu\nu K}T_{\rho\nu}^{acK}e_{\mu c}e^{\rho b}+T_{ab}^{\mu\nu K}T_{\rho\sigma}^{cdK}e_{\mu c}e_{\nu d}e^{\rho a}e^{\sigma b}\right)$
after choosing $\lambda_{I}=-2/3$ so that the vector and axial sectors have canonical normalization. A further rescaling of tensor fields and of the coupling,
$T_{ab}^{\mu\nu K}\rightarrow\sqrt{8/3}T_{ab}^{\mu\nu K},\text{ }g\rightarrow\sqrt{3/8}g$ (45)
puts the tensor sector in canonical form as well. The result is a fully unified quadratic Lagrangian in which a single coupling $g$ controls all gauge submultiplets.
4 Tetrad Condensation – Lifting non-compact directions
The full $SL(2N,C)$ gauge theory contains many more degrees of freedom than are needed to describe known low-energy physics. Our strategy is to show that a dynamical tetrad, subject to the length constraint, can spontaneously break $SL(2N,C)$ and remove the unwanted non-compact directions. In this picture the tetrad not only defines the spacetime geometry but also selects the internal symmetry and filters the spin-connection multiplets that remain light. The low-energy theory then exhibits only $SL(2,C)\times SU(N)$, with $SU(N)$ vectors and a single tensor singlet, while hyperflavored axial and tensor fields are lifted to high mass.
4.1 Dynamical tetrads
We now promote the tetrad to a dynamical field and assign it a gauge-covariant kinetic term of canonical form,
[…]
where the covariant derivative is built from the full spin-connection multiplet (23). Once such a term is added, the theory automatically extends from pure $SL(2N,C)$ to its inhomogeneous version $ISL(2N,C)$, containing local hypertranslations generated by $P_{aK}$. The tetrads $e_{\mu}^{\;aK}$ now act as the corresponding gauge fields.
After rescaling the tetrad to a dimensionless field,
$e_{\mu}^{\;aK}=M_{\!}\,E_{\mu}^{\;aK}$ (47)
where $M$ is a characteristic mass scale, the nonlinear constraint becomes
$\frac{1}{4}\,e_{\mu}^{\;aK}e^{\mu K}{}_{a}=M_{\!}^{2}$ (48)
analogous to a fixed-length condition in a nonlinear $\sigma$ model. A minimal solution consistent with this constraint is
$\overline{e}_{\mu}^{\;}=M_{\!}\,\delta_{\mu}^{\;a}\,\gamma_{a}\,\lambda^{0},\qquad\lambda^{0}=\sqrt{2/N}\mathbf{1}_{N},$ (49)
which is hyperflavor blind. Since this VEV commutes with $\lambda^{k}$ but not with $\gamma_{5}\lambda^{k}$ or $\gamma_{ab}\lambda^{k}$, the vector fields $V_{\mu}^{k}$ remain massless at this stage, whereas the axial-vector multiplet $A_{\mu}^{k}$ and all tensor fields $T_{\mu}^{abK}$ acquire masses of order $gM$. In terms of group theory, the unbroken symmetry after condensation is $SL(2,C)\times SU(N)$, while the broken generators lie in the coset but still close among themselves to form an $sl(2,C)$ algebra; the Lorentz algebra is therefore realized nonlinearly below the scale $M$.
Consequently the theory exhibits two characteristic mass scales: the Planck scale $M_{P}$ fixing Newton’s constant, and the tetrad-condensation scale $M$ at which $SL(2N,C)$ is broken. Above $M$, spin-connection fields propagate with the Yang–Mills type kinetic terms coming from (3.2); below $M$ one can integrate out the heavy tensor and axial fields. The effective action below $M$ contains
$\frac{M_{P}^{2}}{2}\,R(e)+\frac{1}{4}(De)^{2}+\frac{c_{R}}{M^{2}}\,R^{2}+\ldots$ (50)
where $R(e)$ is the Ricci scalar built from the tetrad and $R^{2}$ denotes the curvature-squared terms. The equations of motion for the heavy spin connection enforce metricity, $D_{\mu}e_{\nu}{}^{a}=0$, up to corrections of order $p^{2}/M^{2}$, so variation with respect to $e_{\mu}{}^{a}$ reproduces Einstein’s equation plus Planck-suppressed higher-curvature and torsion contributions.
The constraint (48) fixes only the overall length of the unified tetrad, leaving a continuous set of possible orientations in hyperflavor space. Moreover, in the generic parameterization (34) the antisymmetric modes $\widehat{e}^{[ab]K}$ are eaten by the tensor and axial multiplets, but symmetric hyperflavored modes $\widetilde{e}^{\{ab\}k}$ remain as physical fields and must be given large masses. Both issues are resolved once quantum fluctuations of the gauge fields are included: they generate a Coleman–Weinberg (CW) potential [16] that selects the hyperflavor-blind vacuum and lifts the hyperflavored symmetric modes.
4.2 Hyperflavor-blind vacuum
The length constraint may be imposed via a Lagrange multiplier,
[…]
which fixes the norm but leaves the internal orientation undetermined. A convenient parametrization of the vacuum is
$\bar{e}_{\mu}\;=\;v_{\mu}\otimes H,\qquad v_{\mu}=M\,\delta_{\mu}^{\,a}\gamma_{a},\qquad H^{\dagger}=H,$ (52)
so that the hyperflavor dependence resides entirely in the Hermitian matrix $H$. The choice $H\propto\mathbf{1}_{N}$ realizes a hyperflavor-blind vacuum (HBV) that preserves the residual $SU(N)$ and evidently satisfies (51). It is, however, necessary to show that quantum corrections select this orientation as the true minimum.
The one-loop CW potential for the vector fields in a background $\bar{e}$ is
$U_{V}(e)\;=\;\frac{3}{64\pi^{2}}\sum_{n}m_{n}^{4}(e)\left(\ln\frac{m_{n}^{2}(e)}{\mu^{2}}-\frac{5}{6}\right),$ (53)
where $m_{n}^{2}$ are eigenvalues of the vector mass matrix $M_{V}^{2}(e)$ and $\mu$ is the renormalization scale. For $SU(N)$ generators $T_{V}=\lambda_{k}$ one finds
$(M_{V}^{2})_{kl}(\bar{e})\;$ $=$ $\;g^{2}\mathrm{Tr}\!\left([\lambda_{k},\bar{e}_{\mu}][\lambda_{l},\bar{e}^{\mu}]\right)$
$=$ $4g^{2}M^{2}\,\mathrm{Tr}\!\left([\!\lambda_{k},H][\!\lambda_{l},H]\right)$
where we used
$\lambda_{k},\bar{e}_{\mu}]\;=\;v_{\mu}\otimes[\,\lambda_{k},H\,]\,.$ (55)
If $H$ is proportional to the identity, all commutators vanish, the vector mass eigenvalues are zero, and $U_{V}$ vanishes. For any other orientation some $m_{n}^{2}$ become positive and the CW potential rises above zero. Thus the vector sector uniquely selects the hyperflavor-blind direction as the absolute minimum.
The axial and tensor contributions to the CW potential, $U_{A}(e)$ and $U_{T}(e)$, are insensitive to small deviations in $H$ near the HBV: their masses are set by $gM$ and depend primarily on the Dirac structure, so they contribute only an additive constant at $H\propto\mathbf{1}_{N}$. They do, however, influence the curvature of the potential along directions parametrized by symmetric tetrad fluctuations and thereby generate masses for hyperflavor-carrying symmetric modes.
Counting tetrad degrees of freedom at HBV proceeds as follows. A real $4\times 4$ matrix carries $16$ components, which decompose into $6$ antisymmetric and $10$ symmetric ones per internal index $K$, giving $16N^{2}$ components overall. At the HBV, the antisymmetric $6N^{2}$ modes (linked to torsion) are absorbed by the tensor gauge fields $T_{\mu}^{abK}$. Among the symmetric $10N^{2}$ modes, the $(N^{2}-1)$ adjoint traces $\widetilde{e}_{\mu}^{\mu k}$ become longitudinal for the axial vectors $A_{\mu}^{k}$. This leaves $9N^{2}+1$ symmetric traceless modes: $9(N^{2}-1)$ in the $SU(N)$ adjoint and $10$ in the singlet sector, of which $9$ are graviton-like spin-2 modes and one corresponds to the trace fixed by the length constraint.
Near the HBV, the axial and tensor loops generate a positive-definite CW mass matrix along the hyperflavor-adjoint symmetric directions. Schematically,
$M_{A,T}^{2}(e)\;\propto\;g^{2}\,\mathrm{Tr}\!([T_{A,T},e_{\mu}][T_{A,T},e^{\mu}]),\qquad T_{A}=\gamma_{5}\lambda^{k},\;\;T_{T}=\gamma_{ab}\lambda^{K}$ (56)
and expanding $e_{\mu}=\bar{e}_{\mu}+\varepsilon_{\mu}$ around the HBV, one obtains at quadratic order
$\Delta\mathcal{L}_{\mathrm{CW}}^{(2)}\;=\;-\frac{1}{2}\varepsilon^{k}\cdot\mathcal{M}^{2}\cdot\varepsilon^{k},\qquad\mathcal{M}^{2}\sim\frac{g^{4}}{16\pi^{2}}\,M^{2}$ (57)
for the $9(N^{2}-1)$ adjoint symmetric modes. For the $SU(N)$-singlet graviton fluctuation $\varepsilon^{0}\propto\lambda^{0}$, one has
$\bar{e}_{\mu},\,\varepsilon^{0}]=0\,$ (58)
so neither tree-level mixing with heavy fields nor axial/tensor loops give it a mass. Thus, as required, the gravitational sector is described entirely by the flavor-singlet tetrad,
$e_{\mu}^{\;aK}(x)\;\longrightarrow\;e_{\mu}^{\;a}(x)\,\delta^{K0},\text{ }e_{\mu}^{\;a}=\varepsilon_{\mu}^{a0}$ (59)
while tetrad components with $K\neq 0$ either are eaten by gauge fields (Lorentz-antisymmetric modes) or acquire large CW masses (symmetric adjoints) and thus decouple.
Summarizing this stage: (i) the dynamical tetrad VEV breaks the non-compact directions of $SL(2N,C)$ while preserving a compact hyperflavor $SU(N)$; (ii) local Lorentz invariance remains, realized nonlinearly; (iii) the theory contains two scales, $M_{P}$ and $M$, which need not coincide; and (iv) the low-energy theory recovers Einstein–Cartan gravity with corrections suppressed by $1/M^{2}$ and torsion effects governed by the heavy spin connection. The tetrad-condensation mechanism thus provides a controlled way of filtering the gauge multiplet down to the physically relevant degrees of freedom.
5 Emerging gravity
5.1 Preliminaries
We now turn to the emergence of the Einstein–Cartan term from radiative corrections. The hyperunified $SL(2N,C)$ gauge theory described above is formulated in a Yang–Mills-like fashion, with all field strengths entering quadratically and the tetrad treated as a dynamical field. Its invertibility is enforced through the nonlinear constraint (48), which triggers spontaneous breaking of $SL(2N,C)$ to $SL(2,C)\times SU(N)$ via the hyperflavor-blind vacuum (49) and reduces the tetrad to the neutral component (59). After these steps the low-energy gauge multiplet consists of the $SU(N)$ vectors and the flavor-singlet tensor connection $T_{\mu}^{ab}$.
At this stage the Lagrangian need not contain a tree-level Einstein–Cartan term linear in $T$, even though such a term is allowed by symmetry and by the presence of tetrads. Instead, we will show that the linear curvature term is naturally generated at one loop by fermion bubbles that involve both the tensor connection and the tetrad. Two distinct but complementary routes may be considered, depending on the masses of the fermions contributing to the loops. In a “threshold” scenario, heavy vectorlike fermions of mass $m_{\psi}$ induce a finite, regulator-independent Einstein–Cartan term proportional to $N_{f}m_{\psi}^{2}$ (for $N_{f}$ species), reminiscent of Sakharov’s induced gravity [17] and related effective field theory analyses [18, 19, 20]. In an “universal cutoff” scenario, if gravity or some UV completion provides a physical cutoff $\Lambda\sim M_{Pl}$ for all loop integrals, even light fermions induce an Einstein–Cartan term of order $N_{f}\Lambda^{2}$; this is regulator-dependent but fixes the correct overall scale. In both pictures the Einstein–Cartan term becomes radiative in origin, expressed in terms of the underlying $SL(2N,C)$ parameters while maintaining the unified structure of the quadratic sector.
5.2 Radiative corrections
After tetrad condensation, the quadratic curvature sector is governed by the ghost-free Lagrangian $\mathcal{L}_{G}^{(2)}$ (38) embedded into the $SL(2N,C)$ framework. A separate linear curvature term of the Einstein–Cartan form for the contributing tensor field submultiplet
$\mathcal{L}_{\mathrm{EC}}=\frac{1}{2\kappa}\,e\,e_{[a}{}^{\mu}e_{b]}{}^{\nu}R^{ab}{}_{\mu\nu}[T$], (60)
is allowed by symmetry but need not appear at tree level. We show that this structure with curvature $R^{ab}{}_{\mu\nu}[T]=T^{ab}{}_{\mu\nu}$ is generated radiatively by fermion loops in which both the tetrad and the tensor field couple to matter.
The relevant interaction vertices follow from the fermion kinetic term and the $SL(2N,C)$ covariant derivative (25) when specialized to the neutral tensor field. The basic insertions in the fermion loops are
$V_{e}$ $:\ \frac{i}{M}\,e^{\mu}{}_{a}\,\bar{\Psi}\,\gamma^{a}\overleftrightarrow{\partial_{\mu}}\Psi,$
$V_{T}$ $:\ \frac{i\,g}{4M}\,e^{\mu}{}_{c}\,\bar{\Psi}\,\gamma^{c}\gamma_{ab}\,\Psi\,T_{\mu}^{ab}\,$, (61)
where we have used the rescaling (47) and $e^{\mu}{}_{a}\sim M$. Each vertex carries a factor $1/M$, with $V_{e}$ bringing an external momentum and $V_{T}$ containing the coupling $g$. Because of the explicit $e^{\mu}{}_{c}$ factor in $V_{T}$, each insertion involves both an external $T$ leg and an external tetrad leg.
The minimal one-loop origin of the operator $e\wedge e\wedge R[T]$ is then a pair of fermion-bubble diagrams containing either one $V_{T}$ and one $V_{e}$ insertion ($V_{T}\times V_{e}$) or two $V_{T}$ insertions ($V_{T}\times V_{T}$). The first diagram produces a three-point function corresponding to $e\wedge e\wedge dT$, while the second produces $e\wedge e\wedge(g\,T\wedge T)$. Taken together they reconstruct the full $e\wedge e\wedge R[T]$ structure with $R[T]=dT+gT\wedge T$.
TTeeee
TTeeTTee
Figure 1: Fermion-bubble diagrams generating the Einstein–Cartan term. Left: a mixed $V_{T}$$V_{e}$ insertion produces $e\wedge e\wedge dT$. Right: two $V_{T}$$V_{e}$ insertions produce $e\wedge e\wedge(gT\wedge T)$. Their sum reconstructs $e\wedge e\wedge R[T]$.
The coefficient of (60) depends on the fermion content and on the UV completion. Two complementary scenarios are illustrative. In a “threshold” scenario, vectorlike fermions with masses $m_{\psi}\sim M_{Pl}$ generate a regulator-independent contribution
$\mathcal{L}_{\mathrm{EC}}^{(t)}\sim\frac{N_{f}}{16\pi^{2}}\,m_{\psi}^{2}\,e\wedge e\wedge R[T],$ (62)
so that $M_{Pl}^{2}\sim(N_{f}/16\pi^{2})m_{\psi}^{2}$. In a “universal cutoff” scenario, quantum gravity itself provides a physical cutoff $\Lambda\sim M_{Pl}$, and even light (or massless) fermions induce
$\mathcal{L}_{\mathrm{EC}}^{(c)}\sim\frac{N_{f}}{16\pi^{2}}\,\Lambda^{2}\,e\wedge e\wedge R[T],$ (63)
with the Planck scale determined by the UV boundary condition set by the UV completion. In both cases the emergent EC term is expressed in terms of the parameters of the underlying $SL(2N,C)$ gauge theory and its fermion sector. The quadratic curvature terms are already present via the ghost-free Neville-type Lagrangian, so the full low-energy gravitational sector is an Einstein–Cartan-plus-curvature-squared theory with the correct sign and no ghosts.
5.3 Final remarks
These considerations show that the linear curvature term allowed by $SL(2N,C)$ is generically a sum of a bare piece and an induced one-loop contribution. In a threshold scenario, superheavy vectorlike fermions near the Planck scale generate a finite Einstein–Cartan term which may dominate over any tree-level coefficient if the latter is tied to lower mass scales. The Planck mass then becomes essentially radiative in origin, controlled by the multiplicity and masses of heavy fermion states.
In the universal-cutoff picture, quantum gravity or some UV completion supplies a physical scale $\Lambda\sim M_{Pl}$ that renders all radiative corrections finite. The induced Einstein–Cartan term from light or massless fermions then takes the form $\Delta M^{2}\propto N_{f}\Lambda^{2}$, so the observed Planck mass $M_{Pl}^{2}$ can be viewed as the sum $M^{2}+\Delta M^{2}$, where $M^{2}$ is a bare parameter. Assuming $M^{2}$ is tied to much lower scales (e.g. the $SL(2N,C)$ breaking scale or the weak scale) makes the induced contribution dominant.
One may further entertain the idea that the $SL(2N,C)$ theory (including its SM sector) is classically-scale invariant, with all large scales generated radiatively or via the UV completion. In such a setting the Einstein–Cartan term is absent at tree level, so that the Planck scale and the effective EC coupling necessarily arise solely from such radiative corrections. This does not, by itself, solve the hierarchy problem, but it clarifies how a large gravitational scale may arise without introducing Planck-scale VEVs inside the $SL(2N,C)$ theory.
6 From gauge to matter sector – True HUT carriers
6.1 Are quarks and leptons truly elementary?
So far we have concentrated on the gauge sector. We now address the matter sector and its relation to hyperunification. After tetrad condensation the effective local symmetry is $SL(2,C)\times SU(N)$, while the full $SL(2N,C)$ algebra continues to organize multiplets of fields. On the one hand this yields $SL(2,C)$ gauge gravity; on the other, it provides an $SU(N)$ grand unified group. However, because $SL(2N,C)$ multiplets carry both spinor and internal indices, many familiar $SU(N)$ GUT constructions do not fit comfortably, especially when one insists that quarks and leptons be elementary spin-$1/2$ states in fundamental $SL(2N,C)$ representations.
This can be seen already in the simplest case of $SU(5)$, which might be expected to arise from an $SL(10,C)$ hyperunification scheme. Some low-dimensional chiral fermion multiplets (say left-handed) may then be written in $SU(5)\times SL(2,C)$ components as
$\Psi_{L}^{ia}\text{ },\text{\ \ \ }10=(\overline{5},2)$ (64)
and
$\Psi_{L[ai,\text{ }jb]}=\Psi_{L[ij]\{ab\}}+\Psi_{L\{ij\}[ab]},\text{ \ }45=(10,\text{ }3)+(15,\text{ }1)\,.$ (65)
Here $i,j=1,\dots,5$ label $SU(5)$ indices and $a,b=1,2$ label $SL(2,C)$ spinor indices. In these decompositions, antisymmetry in the combined $(ia)$ indices implies antisymmetry in $i,j$ and symmetry in $a,b$, and conversely. One finds that the $\overline{5}$ representation appears naturally, but the usual fermionic decuplet of $SU(5)$ does not emerge from a purely antisymmetric $SL(10,C)$ representation: the tensor (65) instead corresponds to vector and scalar multiplets rather than to a fermionic decuplet. As a result, the standard $SU(5)$ GUT [21], together with many of its supersymmetric [22] or string-inspired extensions [23, 24], cannot be straightforwardly realized in this hyperunified setting.
More generally, GUTs in which quarks and leptons are placed in purely antisymmetric representations (such as $SU(11)$ models [25]) are problematic: the spin content of the resulting states does not match that of observed fermions once the requirement of an underlying $SL(2N,C)$ structure is imposed. To maintain spin-$1/2$ assignments for ordinary quarks and leptons within an $SL(2N,C)$ framework one would need more complicated multiplets involving both upper and lower indices, which tend to be extremely large and introduce many exotic states.
This strongly points toward a composite interpretation: quarks and leptons may emerge as bound states of more fundamental chiral preons. For such preons the $SL(2N,C)$ representation theory becomes much more economical, with preons transforming in fundamental spinor representations and composites living in higher-rank tensors. In the remainder of this section we follow this line of thought and argue that an $SL(16,C)$ hyperunified model with $SU(8)$ hyperflavor provides a natural home for three generations of composite quarks and leptons [10].
6.2 Preons – metaflavors and metacolors
We adopt the preon framework developed in [26]. At very short distances, potentially near the Planck scale, we postulate the existence of $2N$ massless preons, $N$ left-handed and $N$ right-handed, described by independent Weyl spinors $P_{iaL}$ and $Q_{iaR}$ with $i=1,\dots,N$ and $a=1,2$. They carry a common local metaflavor symmetry
$G_{MF}=SL(2N,C)\,$, (66)
acting in the fundamental representation on both $P_{iaL}$ and $Q_{iaR}$. At the fundamental level we take this metaflavor symmetry to be vectorlike, so that $SL(2N,C)$ acts identically on left- and right-handed preons; left–right symmetry is therefore exact in the preon sector. At large distances, the effective symmetry realized on composite states may be reduced to chiral subgroups.
In addition to metaflavor, preons transform under a chiral metacolor gauge group,
$G_{MC}=SO(n)^{L}\times SO(n)^{R}\,$, (67)
in whose vector representation they reside. We denote the corresponding indices by $\alpha$ and $\alpha^{\prime}$ for the left and right metacolor factors, respectively, and write $P_{iaL}^{\alpha}$ and $Q_{iaR}^{\alpha^{\prime}}$ with $\alpha,\alpha^{\prime}=1,\dots,n$. The chiral metacolor interactions confine preons into bound states—quarks, leptons, and possible additional composites—with a characteristic confinement radius
$R_{MC}\sim 1/\Lambda_{MC}\,$,
where $\Lambda_{MC}$ is the scale at which metacolor becomes strong. By assumption, left and right metacolor sectors have comparable confinement scales due to the underlying L–R symmetry.
The choice of orthogonal metacolor has several advantages. Vector representations of $SO(n)$ are free of gauge anomalies, consistent with the preon assignment, and orthogonal groups allow composites in a variety of representations of the metaflavor group $SL(2N,C)$, including tensors with both upper and lower indices. Under the combined symmetry $SL(2N,C)\times SO(n)^{L}\times SO(n)^{R}$ the preons transform as
$P_{iaL}^{\alpha}[N,(1/2,0);(n,1)]\,,\qquad Q_{iaR}^{\alpha^{\prime}}[N,(0,1/2);(1,n)]\,$, (68)
where we have indicated the transformation properties under metaflavor, Lorentz, and metacolor factors.
In the limit where metaflavor gauge interactions are formally turned off, the preon system exhibits an additional global chiral symmetry,
$K(N)=SU(N)_{L}\times SU(N)_{R}\,$, (69)
acting independently on left- and right-handed preons. A bilinear condensate $\langle P_{L}Q_{R}\rangle$ that would generate large Dirac masses of order $\Lambda_{MC}$ for composites is forbidden by this chiral symmetry. This is a necessary (though not sufficient) condition for the existence of massless composite fermions. In fact, massless composites are expected only when the preon chiral symmetry (69) survives at large distances, and the corresponding pattern is constrained by the ’t Hooft anomaly matching (AM) condition [27]. Roughly speaking, the $SU(N)^{3}_{L,R}$ anomalies carried by preons must match those of the composite fermions in each chiral sector.
An important distinction arises between the vectorlike metaflavor symmetry and the chiral spectator symmetry (69). For the local $SL(2N,C)$, gauge anomalies cancel automatically between left- and right-handed preons and their composites since the theory is vectorlike. For the global chiral symmetry $SU(N)_{L,R}$, however, anomalies must be matched separately in the left and right sectors by suitable combinations of preons and massless composites. In practice one can imagine gauging $SU(N)_{L,R}$ as a spectator symmetry and introducing additional spectator fermions as in [27]. The AM condition then constrains the number of preon species $N$ and the representations in which the massless composite states transform.
6.3 Composites – $SL(16,C)$ and residual $SL(2,C)\times SU(8)$
The metaflavor dynamics is governed by the local vectorlike group $SL(2N,C)$, while the chiral symmetry $SU(N)_{L}\times SU(N)_{R}$ remains global. To analyze anomaly constraints one promotes the chiral symmetry temporarily to a spectator gauge symmetry with auxiliary gauge fields, allowing one to compute the $SU(N)_{L}^{3}$ and $SU(N)_{R}^{3}$ anomalies produced by preons and composites. In this auxiliary theory anomalies in the $SU(N)_{L}$ and $SU(N)_{R}$ sectors are cancelled by suitable spectra of spectator fermions, but the AM condition enforces equality of the anomalies carried by preons and by composites in each sector separately.
In the $SL(2N,C)$ metaflavor theory, the cubic $SL(2N,C)^{3}$ anomalies cancel between left- and right-handed states for any $N$, since the theory is vectorlike. The nontrivial constraint comes instead from the chiral symmetry (69). Let $a(N)$ be the anomaly coefficient for $SU(N)$ in the fundamental representation (so $a(N)=\pm 1$ for left- and right-handed preons), and let $a(r)$ denote the anomaly coefficient for a given composite representation $r$. If $i_{r}$ counts the net number of massless composites in representation $r$ (positive for left-handed, negative for right-handed states), the AM condition takes the form
$n\,a(N)=\sum_{r}i_{r}\,a(r)\,$, (70)
where $n$ is the number of preons of a given chirality. The coefficients $a(r)$ depend on $N$, so (70) can in principle fix $N$ once a particular pattern of composite representations is specified.
To sharpen the constraints we impose three assumptions: (i) all massless composites contributing to (70) are spin-$1/2$; (ii) only the simplest three-preon composites formed by metacolor interactions are present; (iii) all massless composites form a single irreducible representation of the hyperunified symmetry, which at low energies reduces to a single irreducible representation of $SL(2,C)\times SU(N)$. Assumption (i) means that higher-spin components of generic $SL(2N,C)$ multiplets must become massive and decouple, which in turn implies that the preon metaflavor symmetry is effectively reduced to $SL(2,C)\times SU(N)$ at large distances, consistent with the gauge-sector analysis of Secs. 2–4. Assumptions (ii) and (iii) significantly restrict the possible composite spectra.
Under these assumptions, only three-preon composites built from $P_{L}$ (and analogously from $Q_{R}$) are considered. The possible $SU(N)$ representations for spin-$1/2$ composites are third-rank tensors of the form
$\Psi_{L,R}^{\{ijk\}}\,,\quad\Psi_{[ijk]L,R}\,,\quad\Psi_{\{[ij]k\}L,R}\,,\quad\Psi_{\{jk\}L,R}^{i}\,,\quad\Psi_{[jk]L,R}^{i}\,$, (71)
with symmetrization or antisymmetrization in the flavor indices $i,j,k=1,\dots,N$. Evaluating the anomaly coefficients $a(r)$ for these representations and imposing (70) with $n=3$, $a(N)=1$, and $i_{r}=1$ for a single composite representation $r=r_{0}$ yields
$3=a(r_{0})\,.$ (72)
One finds that an integer solution for $N$ exists only when $r_{0}$ is the mixed tensor $\Psi_{\,[jk]L}^{i}$ (and its right-handed analogue) and, moreover, only for a specific value of $N$:
$3=\frac{N^{2}}{2}-\frac{7N}{2}-1\,,\qquad N=8\,$, (73)
which corresponds to a $216$-dimensional representation for each chirality. Thus, among all candidate values, only $SU(8)_{L}\times SU(8)_{R}$ can be realized on massless three-preon composites consistent with anomaly matching, pointing uniquely to $SL(16,C)$ as the underlying metaflavor symmetry [^3] Remarkably, due to direct screening effect of the orthogonal metacolor, the AM condition (73) can be extended to the $SO(n)^{L}\times SO(n)^{R}$ metacolor symmetry case by supplementing the above three-preon representation $\Psi_{\,[jk]L}^{i}$ with $p$ composite fundamental one-preon multiplets $\Psi_{iL}$, appropriately screened by metagluons. Thus, the new AM condition acquires the form $n=\frac{N^{2}}{2}-\frac{7N}{2}-1+p$ whose solutions exist only if np=3n-p=3, so the same eightfold chiral symmetry $SU(8)_{L}\times SU(8)_{R}$ is selected even in a general case [26]..
The unique solution of the AM condition with three-preon composites thus picks out $N=8$ and identifies
$SL(16,C)\longrightarrow SL(2,C)\times SU(8)$ (74)
as the relevant metaflavor pattern: the full $SL(16,C)$ is exact at the preon level, whereas the residual $SL(2,C)\times SU(8)$ is the symmetry realized on massless composites at large distances. The corresponding massless multiplets $\Psi_{\,[jk]aL}^{i}$ and $\Psi_{\,[jk]aR}^{\prime\,i}$ fill the $216_{L,R}$ representations of $SU(8)$, with spinor index $a=1,2$.
Upon decomposing the $SU(8)$ multiplet under $SU(5)\times SU(3)$ one obtains (suppressing spinor indices)
$216_{L,R}=[(\overline{5}+10,\text{ }\overline{3})+(45,1)+(5,8+1)+(24,3)+(1,3)+(1,\overline{6})]_{L,R}$ (75)
where $(\overline{5}+10,\overline{3})_{L}$ corresponds precisely to three generations of quarks and leptons forming a triplet of a family symmetry $SU(3)$ [28]. The same pattern appears for $216_{R}$ due to the underlying vectorlike $SL(16,C)$ symmetry. Even though preons remain massless, all composite states listed in (75) are in principle allowed to pair up and acquire heavy masses through Dirac terms.
6.4 L-R symmetry violation – chiral quark-lepton families
To obtain a chiral low-energy spectrum containing three light quark-lepton families, one needs to break the fundamental L-R symmetry of the preon sector so that only the left-handed combination $(\overline{5}+10,\overline{3})_{L}$ remains massless, while its right-handed counterpart acquires a large mass. Following [26], we assume that this breaking is triggered by interactions in the right-handed preon sector, such that the chiral symmetry realized on right-handed composites at large distances is reduced to $SU(5)\times SU(3)]_{R}$, while the left-handed sector retains the full $SU(8)_{L}$.
The corresponding breaking pattern may be caused by some typical L-R symmetric polynomial potential for the third-rank antisymmetric scalars $\Phi_{[pqr]}$ and $\Phi_{[pqr]}^{\prime}$ composed from $P$ and $Q$ preons, respectively. This potential, after breaking of $SL(16,C)$ to $SL(2,C)\times SU(8)$, can be written in the form
$U=-M_{U}^{2}(\Phi^{2}+\Phi^{\prime 2})+h_{1}(\Phi^{2}\Phi^{2}+\Phi^{\prime 2}\Phi^{\prime 2})^{2}+h_{2}(\Phi^{4}+\Phi^{\prime 4})+h_{3}(\Phi^{2}\Phi^{\prime 2})$ (76)
which presumably are induced by the multi-preon interactions at large distances (here the notations $\Phi^{2}=Tr(\Phi^{+}\Phi)$, $\Phi^{\prime 2}=Tr(\Phi^{\prime+}\Phi^{\prime})$, $\Phi^{4}=Tr(\Phi^{+}\Phi\Phi^{+}\Phi)$ and $\Phi^{\prime 4}=Tr(\Phi^{\prime+}\Phi^{\prime}\Phi^{\prime+}\Phi^{\prime})$ were used). Likewise, these interactions may also produce, among others, the invariant Yukawa-type couplings for the composite multiplets (75)
$L_{Y}=\frac{1}{M_{Y}^{2}}\left(\overline{\Psi}_{Lr}^{[pq]}\widehat{D}\Psi_{L[tu]}^{s}\Phi_{[pqs]}\Phi^{[rtu]}+\overline{\Psi}_{Rr}^{[pq]}\widehat{D}\Psi_{R[tu]}^{s}\Phi_{[pqs]}^{\prime}\Phi^{\prime[rtu]}\right)$ (77)
(p,q,r,s,t,u=1,,8p,q,r,s,t,u=1,...,8) being initially L-R symmetric as well. The mass parameter $M_{Y}$ is some UV cutoff which in our case can be ultimately related to the preon confinement energy scale $\Lambda_{MC}$.
As in the known left-right models [29], in the potential (76) for a natural range of the parameters, particularly, for $h_{3}>2(h_{1}+h_{2})$, and properly chosen the higher dimension coupling constants, the scalars $\Phi_{[ijk]}$ and $\Phi_{[ijk]}^{\prime}$ may develop the totally asymmetric VEV configuration
$\left\langle\Phi^{[pqr]}\right\rangle=0\text{ , \ }\left\langle\Phi^{\prime[pqr]}\right\rangle=\delta_{\overline{p}}^{p}\delta_{\overline{q}}^{q}\delta_{\overline{r}}^{r}\epsilon^{\overline{p}\overline{q}\overline{r}}M_{LR}\text{ \ \ \ }(\overline{p},\overline{q},\overline{r}=1,2,3)$ (78)
which leads at the scale $M_{LR}$ to the spontaneous breakdown of the starting $SU(8)$ down to the $SU(5)\times SU(3)$ symmetry group which we may identify as the conventional $SU(5)$ grand unification times the $SU(3)_{F}$ family symmetry. Moreover, due to asymmetric VEV (78), the Yukawa couplings (77) certainly lead to the chiral symmetry breaking for the right-handed composites, $SU(8)_{R}\rightarrow[SU(5)\times SU(3)]_{R}$, leaving intact the $SU(8)_{L}$ symmetry for the left-handed ones. In terms of the spectator gauge chiral symmetry, this means that all non-diagonal gauge bosons related to the broken generators of the coset $SU(8)_{R}\ /$ $SU(5)\times SU(3)]_{R}$ acquire at large distances the scale $M_{LR}$ order masses. This means that, though the massless right-handed preons still possess the $SU(8)_{R}$ symmetry, the masslessness of their composites at large distances is now solely controlled by its remained $SU(5)\times SU(3)]_{R}$ part. Indeed, the right-handed composites no longer fill the complete $216_{R}$ multiplet but instead satisfy a modified anomaly matching condition based on $SU(5)\times SU(3)]_{R}$. One finds that the right-handed spectrum consistent with these anomalies excludes the multiplet $(5+10,3)_{R}$, just leaving the submultiplet combination
$(45,1)+(5,8+1)+(5+\overline{5},3)+(1,\overline{6})]_{R}$ (79)
which satisfies the $SU(5)^{3}$ and $SU(3)^{3}$ anomaly matching conditions with the right-handed $Q$ preons. When the left-handed multiplet (75) and the reduced right-handed multiplet (79) are combined, all components except $(\overline{5}+10,\overline{3})_{L}$ can pair up and become heavy, leaving only
[…]
as the anomaly-free light spectrum, with $(24+1,3)_{L}$ and $(5+\overline{5},3)_{R}$ expected to be lifted to higher scales through the family-symmetry interactions. Thus three chiral families of quarks and leptons naturally remain light, transforming as $(\overline{5}+10,\overline{3})_{L}$.
Once L-R symmetry is broken in this way, the initially vectorlike metaflavor symmetry $SU(8)\times SL(2,C)$ effectively reduces in the composite sector to a chiral subgroup $SU(5)\times SU(3)_{F}]\times SL(2,C)$, where $SU(3)_{F}$ is a family symmetry acting on the three quark-lepton generations [28]. Further symmetry breaking is then driven by scalar multiplets (which may themselves be composite) that break $SU(5)$ and $SU(3)_{F}$ to the Standard Model, generating the observed pattern of gauge bosons and fermion masses and mixings.
6.5 Down to Standard Model
A conventional scenario for further breaking of the $SU(5)\times SU(3)_{F}$ invariance in the theory—considered as the appropriate GUT and family symmetry candidates—introduces a suitable set of scalar fields that may reduce this theory down to the Standard Model. For this, one takes the $SL(16,C)$ adjoint scalar multiplets of the type
$\Sigma=(\sigma^{k}+i\sigma_{5}^{k}\gamma_{5})\lambda^{k}+\sigma_{ab}^{K}\gamma^{ab}\lambda^{K}/2$ (81)
(k=1,,64;k=1,...,64; $K=0,k$) which transform as
$\Sigma\rightarrow\Omega\Sigma\Omega^{-1}\text{\ }$ (82)
In general, $\Sigma$ contains not only scalar components but also pseudoscalar and spinorial tensor components in (81). These can be filtered out by the covariant tetrad-constraint mechanism. Imposing on $\Sigma$ the constraint
$\Sigma=e_{\rho}\Sigma e^{\rho}/4$ (83)
one finds, after tetrad condensation, that $\Sigma$ is effectively “sandwiched” by the neutral tetrads (32). As a result, only the pure scalar components remain in the $SU(5)$ symmetry breaking multiplet, $\Sigma=\sigma^{k}\lambda^{k}$, thereby enabling the breaking of the $SU(5)$ GUT down to the Standard Model, $SU(5)\rightarrow SU(3)_{c}\times SU(2)_{W}\times U(1)_{Y}$, caused by the conventional potential [29].
As to the final breaking of the SM and accompanying family symmetry $SU(3)_{F}$ to $SU(3)_{c}\times U(1)_{em}$, it occurs through the extra multiplets with assignments determined by representations chosen for quarks and leptons. They are $H^{[pa,qb,rc,sd]}$, and $\xi_{[pa,qb]}$ and $\chi_{\{pa,qb\}}$ of $SL(16,C)$, which contain an even number of antisymmetrized and symmetrized $SU(8)$ and $SL(2,C)$ indices (p,q,r,s,t,u=1,,8;p,q,r,s,t,u=1,...,8; $a,b,c,d,e,f=1,2$). These multiplets include, among others, the genuine scalar components which develop the corresponding VEVs giving masses to the weak bosons, as well as to the family bosons of the $SU(3)_{F}$. They also generate masses for quarks and leptons located in the left-handed fermion multiplet (75) through the $SL(16,C)$ invariant Yukawa couplings. Particularly, for the submultiplet $(\overline{5}+10,\overline{3})_{L}$ one has, as in a conventional $SU(5)$ GUT, the two independent couplings
$\frac{1}{\emph{M}}\left[\Psi_{[qa,\text{ }rb]L}^{pc}C\Psi_{[te,\text{ }uf]L}^{sd}\right]H^{\{[qa,rb],[te,uf]\}}\left(a_{U}\xi_{[pc,sd]}+b_{U}\chi_{\{pc,sd\}}\right),$
$\frac{1}{\emph{M}}\left[\Psi_{[qa,\text{ }rb]L}^{pc}C\Psi_{[pc,\text{ }te]L}^{sd}\right]H^{\{[qa,rb],[te,uf]\}}(a_{D}\xi_{[sd,uf]}+b_{D}\chi_{\{sd,uf\}})$ (84)
with distinct index contraction for the up quarks, and down quarks and leptons, respectively. The mass $M$ represents an effective scale in the theory that, in the composite model of quarks and leptons, can be linked to their compositeness scale $\Lambda_{MC}$, while $a_{U,D}$ and $b_{U,D}$ are dimensionless constants of the order of $1$.
Actually, these couplings contain two types of scalar-containing multiplets with the following $SU(8)\times SL(2,C)$ components—the $H$ multiplet with the scalar components
$H^{[pqrs]\{[ab],[cd]\}}(70,1)$ (85)
and the $\xi$ and $\chi$ multiplets, whose scalar components look as
$\xi_{[pq\boldsymbol{]}[ab]}(28,1),\text{ \ }\chi_{\{pq\boldsymbol{\}}[ab]}(36,1)$ (86)
Decomposing them into the components of the final $SU(5)\times SU(3)_{F}$ symmetry one finds the full set of scalars
$70$ $=$ $(5,1)+(\overline{5},1)+(10,\overline{3})+(\overline{10},3)$
$28$ $=$ $(5,3)+(10,1)+(1,\overline{3})$
$36$ $=$ $(5,3)+(15,1)+(1,6)$ (87)
containing the $SU(5)$ quintets $(5,1)$ and $(\overline{5},1)$ to break the Standard Model at the electroweak scale $M_{SM}$, and the $SU(3)_{F}$ triplet and sextet, $(1,\overline{3})$ and $(1,6)$, to properly break the family symmetry at some larger scale $M_{F}$. One may refer to the scalars (85) and (86) as the respective ”vertical” and ”horizontal” scalars, which provide the simplest form of the above Yukawa couplings. Acting in pairs, they presumably determine the masses and mixings of all quarks and leptons. Lastly, and importantly, in the model under consideration, these scalars may themselves be composite states formed from the same preons as quarks and leptons [26].
After the $SU(8)$ symmetry breaking the Yukawa couplings (84) acquire the transparent $SU(5)\times SU(3)_{F}$ invariant form (all metaflavor indices are omitted)
[…]
$[(\overline{5},\overline{3})_{L}C(10,\overline{3})_{L}](\overline{5},1)_{D}[a_{D}(1,\overline{3})+b_{D}(1,6)]/\mathcal{M}\text{ }$ (88)
where we only include those components of the vertical scalar $H$ and horizontal scalars $\xi$ and $\chi$ which develop the VEVs. Just the horizontal scalar VEVs determine through the Yukawa couplings (88) the mass matrices for quarks and leptons
$\widehat{m}_{\overline{p}\overline{q}}^{U}$ $=$ $\left\langle 5,1\right\rangle_{U}[a_{U}\left\langle 1,\overline{3}\right\rangle_{[\overline{p}\overline{q}]}+b_{U}\left\langle 1,6\right\rangle_{\{\overline{p}\overline{q}\}}]/\mathcal{M}\text{ \ }$
$\widehat{m}_{\overline{p}\overline{q}}^{D}$ $=$ $\left\langle\overline{5},1\right\rangle_{D}[a_{D}\left\langle 1,\overline{3}\right\rangle_{[\overline{p}\overline{q}]}+b_{D}\left\langle 1,6\right\rangle_{\{\overline{p}\overline{q}\}}]/\mathcal{M}\text{ \ }$ (89)
where the angle brackets denote the corresponding VEVs, while $\overline{p},\overline{q}=1,2,3$ stand for family indices. The matrices $\widehat{m}_{\overline{p}\overline{q}}^{U}$ and $\widehat{m}_{\overline{p}\overline{q}}^{D}$ are defined at the grand unified scale and have to be then extrapolated down to an actual mass range for quarks and leptons. Depending on which components of the above symmetrical and asymmetrical VEVs are developed, one comes to different texture zero types for all matrices involved. The strong hierarchies of the quark-lepton masses and mixings may be now explained by somewhat softer hierarchies between the breaking directions of the $SU(3)_{F}$ family symmetry whose scale $M_{F}$ is imposed to be close to the effective scale $M$
The composite nature of quarks and leptons introduces, in addition to the Planck scale $M_{P}$ and the $SL(16,C)$ breaking scale $M$ (determined by the tetrad condensate), a new fundamental scale: the preon confinement scale $\Lambda_{MC}$. The phenomenology depends on the relative hierarchy among these three scales. If $M$, $\Lambda_{MC}$ and $M_{P}$ are all near one another, the framework mainly provides a structural reinterpretation of known physics. If $M$ and $\Lambda_{MC}$ are considerably below $M_{P}$, a rich spectrum of preonic and hyperunified states may become accessible, offering possible experimental windows into the underlying $SL(16,C)$ hyperunification.
7 Summary and outlook
[…]
The low-energy symmetry that remains after tetrad condensation is $SL(2,C)\times SU(N)$, realized on a spectrum in which only the neutral tetrad (graviton) and the $SU(N)$ vector multiplet are massless. Starting from a unified quadratic-strength Lagrangian with a single gauge coupling, we argued that the Einstein–Cartan term can be induced by fermion loops involving the flavor-singlet tensor connection and the tetrad, with either heavy vectorlike fermions near the Planck scale or a universal UV cutoff providing the required scale.
On the matter side, we saw that attempting to assign elementary quarks and leptons to $SL(2N,C)$ multiplets usually leads to large representations with unwanted spin content. This motivates a composite viewpoint in which quarks and leptons are bound states of preons in fundamental $SL(2N,C)$ representations. Applying the ’t Hooft anomaly matching condition to the global chiral symmetry $SU(N)_{L}\times SU(N)_{R}$ of preons, while assuming that only three-preon spin-$1/2$ composites remain massless and fit into a single representation, singles out $SU(8)_{L}\times SU(8)_{R}$ at large distances and thus $SL(16,C)$ as the hyperunified metaflavor group. The residual symmetry $SL(2,C)\times SU(8)$ naturally yields three families of composite quarks and leptons, organized into an $SU(5)\times SU(3)_{F}$ structure after L–R breaking in the composite sector.
The proposed hyperunification mechanism is purely four-dimensional and gauge-based; it does not rely on extra dimensions [30, 31, 32] or string excitations [33]. The tetrad, which is required in any case to couple spinors to gravity, is used as the field that condenses, fixes the vacuum, and lifts non-compact directions with a clear understanding of which gauge components have been decoupled.
Further work should complete the preon model for $N=8$, including the full set of composites. If the tetrad-condensation scale $M$ and the preon confinement scale $\Lambda_{MC}$ lie well below $M_{Pl}$, remnants of the filtered hyperunified structure could leave characteristic signatures in the spectrum of heavy bosonic and fermionic composites. A related possibility is that all $SL(2N,C)$ gauge fields themselves may be composite, built from preon bilinears with only global $SL(2N,C)$ imposed at the fundamental level [34, 35, 36, 37, 38, 39]. Extending the present work to this fully composite setting remains an interesting direction for future study.
Acknowledgments
I am grateful to Colin Froggatt, Oleg Kancheli, Holger Nielsen and Grigory Volovik for illuminating discussions, and Lin Adams, George Chkareuli and Igor Gelbach for reading the manuscript and useful remarks.